OpenStudy (hkmiu):

Find the missing value when given the modulus. |a-i| = sqrt37 (in the third quadrant)

OpenStudy (anonymous):

\[|a + bi| = \sqrt {a^2 + b^2}\]so\[\left| a-i \right|=\sqrt{37} \implies a= \pm 6\]The third quadrant means a = -6.

OpenStudy (hkmiu):

Okay thank you, just making sure!

OpenStudy (anonymous):

No sweat. Do math every day.